OpenStudy (kayders1997):

Ok if something says d/dx than integral (x^2+4)^5 is the answer just what is already shown?

jimthompson5910 (jim_thompson5910):

Yes, \[\Large \frac{d}{dx}\int(x^2+4)^5dx = (x^2+4)^5\] since, \[\Large \frac{d}{dx}\int f \ '(x)dx = f(x)\]

OpenStudy (kayders1997):

Thank you @jim_thompson5910 :)

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (kayders1997):

@jim_thompson5910 it wouldn't matter if the d/dx was inside the integral right?

jimthompson5910 (jim_thompson5910):

if it's inside the integral, then you'll have \[\Large \int \frac{d}{dx}\left(f(x)\right)dx = f(x)+C\] where C is a constant

OpenStudy (kayders1997):

Oh so your just adding a c? so it would be (x^2+4)^5+c

jimthompson5910 (jim_thompson5910):

yes, usually it's capital C

OpenStudy (kayders1997):

oh ok

jimthompson5910 (jim_thompson5910):

\[\Large \int f \ '(x)dx = f(x)+C\]

jimthompson5910 (jim_thompson5910):

\[\Large \frac{d}{dx}\Big[f(x)\Big] = f \ '(x)\]

OpenStudy (kayders1997):

ok, yes that makes sense